The pH of an aqueous solution of 0.456 M trimethylamine (a weak base with the formula (CH3)3N) is
Answer #1

The ionization of  (CH3)3N) in water is represented as

(CH3)3N (aq)+ H2O(l) → (CH3)3NH+(aq) + OH (aq)

As  (CH3)3N is a weak base and ionizes to a lesser extent, we must use Kb value to determine equilibrium concentration of OH

Setting up ICE table to determine equilibrium concentration of all species

(CH3)3N (aq) (CH3)3NH+(aq) OH (aq)
Initial conc 0.456 M 0 0
change in conc -x M +x M +x M
Equilibrium conc (0.456-x) M x M x M

Kb of  (CH3)3N (aq) = [(CH3)3NH+(aq) ][OH (aq)] / [(CH3)3N (aq)]

= x.x / (0.456-x)

=x2/(0.456-x)

Assuming x to be small and ignoring it from denominator

Kb =x2/0.456

Kb of  (CH3)3N (aq) = 6.4 *10-5 -----[literature value]

6.4 *10-5 = x2/0.456

x = 0.0054

Thererefore, at equilibrium [(CH3)3NH+(aq) ]=[OH (aq)] = 0.0054 M

pOH = -log[ OH ] = -log(0.0054)

=2.27

pH+pOH= 14

pH= 14- pOH

= 14 - 2.27

=11.73

pH of 0.456 M CH3)3N = 11.73